DCrandall13:

An equation is shown below: 3(2x – 7) = 3 Part A: How many solutions does this equation have? Part B: What are the solutions to this equation? Show your work.

snowflake0531:

I think that you can do Part B to prove Part A \(The~distributive~property:~a(b-c) = a(b)-a(c)\) so 3(2x)-3(7)=3 Can you solve for x?

CripQUEZZ:

@dcrandall13 u there?